Circular No. 8103

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003cp, 2003cq, 2003cr, 2003cs, 2003ct Additional discoveries of apparent supernovae have been reported from

unfiltered CCD images — SN 2003cp in MCG +10-12-78 by M. Armstrong (cf. IAUC 8074); SN 2003cq in NGC 3978 by R. Arbour (cf. IAUC 8097); SN 2003cr in UGC 9639 by T. Boles (cf. IAUC 8093); and SNe 2003cs and 2003ct by W. M. Wood-Vasey et al. (NEAT images; cf. IAUC 8102).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cp Mar. 30.86 8^{h}09^{m}11.^{s}13 +57^{o}41′14′′.0 16.2 3′′ W, 9′′ N
2003cq Mar. 30.84 11 56 14.17 +60 31 19.1 17.0 32′′.0 E, 2′′.3 S
2003cr Mar. 31.18 14 58 35.53 +44 53 06.4 17.7 4′′.7 W, 5′′.0 N
2003cs Feb. 3.49 11 20 37.52 − 4 54 32.7 17.9 0′′.3 W, 5′′.7 N
2003ct Mar. 27.19 11 22 56.20 − 2 56 59.9 19.0 12′′.5 E, 2′′.9 S

Additional approximate CCD magnitudes (from the respective discoverers unless noted otherwise): SN 2003cp, 2000 Dec. 30 UT, [19.0; 2003 Mar. 14, [18.5; 31.786, 16.2. SN 2003cq, Mar. 14, [17.4; 31.816, 17.9 (Boles); 31.830, 17.0. SN 2003cr, Feb. 24, [19.0; 31.894, 17.5. SN 2003cs, 2002 Apr. 18.14, [20.8; 2003 Feb. 9.44, 17.1; Mar. 27.19, 19.4; 31.22, 19.6. SN 2003ct, Feb. 9.44, [20.6; Mar. 25.20, 19.1; 31.22, 19.1. Nothing is visible at the locations of SNe 2003cp, 2003cq, or 2003cr on Palomar Sky Survey red or blue plates. Arbour adds that a CCD image by D. Briggs on Mar. 30.885 yields position end figures 14^{s}.11, 19′′.3 for SN 2003cq, while Boles measured position end figures 14^{s}.20, 19′′.6 from his own image taken on Mar. 31.816.

SUPERNOVA 2003bw IN IC 1077 Further to IAUC 8102, M. Hamuy reports that a spectrum of SN

2003bw (cf. IAUC 8093), obtained on Mar. 31.39 UT, shows it to be a type-II supernova, at least two months old.

COMET C/2002 W5 (SOHO) Further to IAUC 8100, D. Hammer reports measurements for a Kreutz

sungrazing comet found by S. Hönig from SOHO C2 and C3 coronagraph website images. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2003-F45.

2002 UT  α_{2000}  δ_{2000}
Nov. 19.696  15^{h}27.^{m}1  −23^{o}28′
2003 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green